OpenStudy (anonymous):

If g(x)= 3+x+e^x, find g,inverse,(4)

OpenStudy (anonymous):

In an inverse function, the input or x is the y or output of the uninverted function. So set 3+x+e^x=4. Do you follow

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So you solve, where x+e^x=1.

OpenStudy (anonymous):

got cha

OpenStudy (anonymous):

x=0, right?

OpenStudy (anonymous):

how did you get rid of the e^x

OpenStudy (anonymous):

I didn't really get rid of it, I just realized that e^0=1, which is what I wanted from the problem. You could set the whole thing equal to zero and graph it.

OpenStudy (anonymous):

okay